OpenStudy (anonymous):

what is the derivative of tan(1)?

OpenStudy (yuki):

0

OpenStudy (yuki):

because tan(1) is a constant

OpenStudy (anonymous):

okay then i have to post the actual question bc i am having trouble

OpenStudy (yuki):

okay, shoot it

OpenStudy (anonymous):

use this formula |E_{n}(x)|\le\frac{M}{(n+1)!}|x-a|^{{n+1}} find the bound for the error in approximating the quantity with a third degree polynomial for tan 1, f(x) =tanx about x=0

OpenStudy (yuki):

I have no idea what you plugged in lol are you asking us to find what tan(1) is using the McLaurin series with degree 3 ?

OpenStudy (anonymous):

i have to find the error bound using the lagrange error bound for P(x)

OpenStudy (anonymous):

that the formula that i typed

OpenStudy (yuki):

okay, then let's do it step by step. did you find the Taylor (in our case McLaurin because the center is at x=0) Series of tan (x) ?

OpenStudy (anonymous):

no not yet

OpenStudy (yuki):

to use the Lagrange error bound you need the (n+1)th term in the Taylor polynomial for tan(x). since we are approximating tan(x) by a 3rd degree, we need the 4th term of the Taylor polynomial.

OpenStudy (anonymous):

okay

OpenStudy (yuki):

are you comfortable finding it ?

OpenStudy (anonymous):

i found it i got -6( 1+ x^2)^-4

OpenStudy (anonymous):

oh i did it wrong i wrote down the wrong derivative

OpenStudy (anonymous):

i need helpp

OpenStudy (yuki):

the derivative of tan(x) = sec^2(x) which makes it really hard to find te 4th term. take your time

OpenStudy (anonymous):

i did find it i got -24 sinx/(cosx)^5

OpenStudy (yuki):

now we need to find the upper bound of that term. the problem here is that -24sin(x)/(cosx)^5 has no bound because when cos(x) is close to 0, this number becomes arbitrarily big.

OpenStudy (anonymous):

so what do i do??

OpenStudy (yuki):

give me one second

OpenStudy (anonymous):

ok

OpenStudy (yuki):

were you given any interval ?

OpenStudy (anonymous):

it just says x=0 and tan (1) so 0<x<1

OpenStudy (yuki):

nath lol that info was actually very important lmao

OpenStudy (yuki):

since x is between 0 and 1, we just need to find the upper bound of -24 sin/cos^5 between those intervals

OpenStudy (yuki):

cosine decreases as x increases from 0 to 1 (radians) and sine increases as x increases from 0 to 1 so the larges number M can be is found by plugging in x =1 into the fourth derivative of tan(x) just to let you know I am not really sure if it is -24sin(x)/cos^5(x), but if you are 100% sure we can proceed

OpenStudy (anonymous):

is there anyway you could do it really quickly just to make sure its correct cause we are make an effort to do the problem

OpenStudy (yuki):

I will try to find it, then. try to trust me for the derivative. even if we don't get the right answer the concept is solid. give me a minute to find f(^4)(x)

OpenStudy (anonymous):

k ty

OpenStudy (yuki):

\[f = tanx\] \[f' = \sec^2x\] \[f" = 2\sec^2x tanx\] \[f^{(3)} = 4secx tanx + 2 \sec^4x\] \[f^{(4)} = 4secx \tan^2x + 4 \sec^3x + 8\sec^4tanx\]

OpenStudy (yuki):

let me see if I can simplify it

OpenStudy (anonymous):

kk

OpenStudy (yuki):

I think this one looks best among others \[f^{(4)} = {2\sin^2(2x) +4\cos^2(x) +8\sin(x) \over \cos^5(x)}\]

OpenStudy (yuki):

so M would be 2sin^2(2)+4cos^2(1) + 8sin(1) over cos(1)

OpenStudy (anonymous):

the thing is i said the derivative of tan x is 1/cos^2 x and i got something completely different from you

OpenStudy (anonymous):

okay lets move on so how do i know what to plug into x

OpenStudy (yuki):

let's try that as well \[f' = { 1 \over \cos^2(x)}\] \[f" = {2\sin(x) \over \cos^4(x)}\] which simplyfies to 2sec^2(x)tan(x) \[f^{(3)} = {2\cos^4(x) +6\cos^2(x)\sin^2(x) \over \cos^6(x)}\] I don't think I really want to do the next one ...

OpenStudy (yuki):

basically your idea is "what is the maximum value of f^(n+1) in the given interval ?" if you are using a Taylor series with center a to apprx, tan(2), the interval is from a < x < 2 if they want you to evaluate tan(x) from -3 < x < 5 for example, you will pick the x that maximizes f^(n+1) between -3 and 5

OpenStudy (anonymous):

so for this problem what my max varies for cos and and sin

OpenStudy (anonymous):

bc their max is different

OpenStudy (yuki):

so if we used the f^(4) I got, what you can say is I cannot calculate sin(1) or cos(1) , but I know that 1 radian is less than 60 degrees and greater than 0. to maximize the expression that I have, I need to make the numerator as small as possible and the denominator as large as possible. so since 0<sin(1)<sin(60), 0<cos(1)<cos(60) I will get the following

OpenStudy (yuki):

ops nvm

OpenStudy (yuki):

i couldn't get a nice number ...

OpenStudy (yuki):

anyway, I don't feel like I was of much help, but the Idea is this. once you get a valid M, you start solving for n. since it will involve factorials, all you have to do is to plug in a couple of integers to see when it gives you a true statement. since tangent is not an easy function to calculate the derivatives, I am not 100% sure with my calculations either.

OpenStudy (anonymous):

alright ty u very much